3A) If both parents are heterozygous carriers of a mutant allele, what is the chance that their first child will be homozygous recessive for the mutation?

3B) Parents who are each heterozygous carriers for a recessive mutant allele have a child who does not have the condition. What is the chance this child is a heterozygous carrier of the condition?

3C) If the first child of parents who are both heterozygous carriers of a recessive mutant allele is homozygous recessive, what is the chance the second child of the couple will be homozygous recessive?

3D) If the first child of parents who are both heterozygous carriers of a recessive mutant allele is homozygous recessive, what is the chance the second child will be a heterozygous carrier of the recessive mutation?
Answer #1

Mm x Mm

M m
M MM Mm
m Mm mm

a. 1/4 or 25% chance of their first child will be homozygous recessive (mm) for the mutation.

b. 2/4 or 1/2 chance that the child is a heterozygous carrier of the condition.

c. Chance of the second child be homozygous recessive is 1/4 x 1/4 = 1/16

d.  Chance the second child will be a heterozygous carrier of the recessive mutation if the first child is homozygous recessive.

1/4 x 1/2 = 1/8

where 1/2 is the chance of being carrier and 1/4 is the chance of being homozygous recessive.